A copper bar is 2.5 m long at 150C. What is the change in the length of the bar when it is heated to 350C?


1
Expert's answer
2022-01-16T13:10:41-0500

The change in length:


"\Delta L=L_0\alpha \Delta T=(2500)(16.7·10^{-6})(35-15)=0.835\text{ mm}."